OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (–2, 6) and has a slope of 4. y – 2 = 4(x + 6) y + 6 = 4(x – 2) y + 2 = 4(x – 6) y – 6 = 4(x + 2)

OpenStudy (karatechopper):

y-6=4(x+2)

OpenStudy (anonymous):

thanks you

hero (hero):

Are you sure that's right?

OpenStudy (anonymous):

wait a second your twelve?

OpenStudy (karatechopper):

yes, and yes hero if u want go ahead and declare me wrong again

OpenStudy (anonymous):

lol r u telling me it's wrong hero?

OpenStudy (karatechopper):

explain it

OpenStudy (karatechopper):

i checked my textbook it is right

OpenStudy (karatechopper):

point slope form is written in y-y1=m(x-x1) form

OpenStudy (anonymous):

if it's in a textbook thatn you know it's right

hero (hero):

Actually, she has it right

hero (hero):

You're lucky

OpenStudy (karatechopper):

exactly and tht was the same way it was for the other question where hero also said i was wrong

OpenStudy (karatechopper):

im not lucky, i just kno my stuff

hero (hero):

You don't know how to graph x = 4 or y = -3

OpenStudy (karatechopper):

yes i do i checked my textbook

hero (hero):

But I'm not here to be your enemy. I'm here to be your friend.

OpenStudy (anonymous):

thats easy even i know that

OpenStudy (karatechopper):

i kno but answer one thing for me: give me the ordered pair for y=-3

hero (hero):

you're still arguing about this? I thought you said you checked your textbook

OpenStudy (karatechopper):

i did but then u said i dont kno how to graph tht so give me the ordered pair or i will regret it for my life

hero (hero):

Angel, post it as a separate question

hero (hero):

I can only see one question per post.

hero (hero):

If I see two questions in the same post, I can comprehend that

hero (hero):

I can't*